Calculate the root mean square velocity of air molecules at N.T.P.(Density of air = 1.293 Kg/m³).


1
Expert's answer
2021-01-10T18:27:35-0500

The root mean square velocity of air molecules can be found from the formula:


"v_{rms}=\sqrt{\dfrac{3RT}{M}}."

Let's express "\dfrac{RT}{M}" in terms of "P" and "\rho". From the ideal gas law, we have:


"PV=nRT,"

here, "n = \dfrac{m}{M}."

Let's divide both sides of the equation by "V":


"P=\dfrac{m}{V}\dfrac{RT}{M}."

Since "\rho=\dfrac{m}{V}" we have:


"P=\rho\dfrac{RT}{M}"

Finally, we get:


"\dfrac{RT}{M}=\dfrac{P}{\rho}."


Then, we can calculate the root mean square velocity of air molecules at normal temperature and pressure:


"v_{rms}=\sqrt{\dfrac{3RT}{M}}=\sqrt{\dfrac{3P}{\rho}},""v_{rms}=\sqrt{\dfrac{3\cdot 1.013\cdot10^5\ Pa}{1.293\ \dfrac{kg}{m^3}}}=485\ \dfrac{m}{s}."

Answer:

"v_{rms}=485\ \dfrac{m}{s}."